OpenStudy (anonymous):

In ∆ABC, if a = 8.75 centimeters, c = 4.26 centimeters, and m B = 87°, what is the length of b to two decimal places?

OpenStudy (anonymous):

9.5cm

OpenStudy (campbell_st):

you will need the law of cosines to solve it \[b^2 = a^2 +c^2 - 2a \times c \times \cos(B)\] just substitute your values